# Time Non-locality in Dark Matter and LSS

### v1

@@ 3 Non-locality in dark matter
SPT has been sucessful in describing LSS at very large scales but as we approach shorter distances i.e. higher wavenumber, linear SPT starts to diverge from the results of fully nonlinear -body simulations. Even for , i.e. quasilinear scales where , higher loop calculations of the SPT power spectrum do not converge to the correct result [32]. To cure this, the correct treatment is to smooth out the equations of motion over a length scale above the non-linear scale. The effect of this smoothing out procedure is to include “effective" contributions from non-linear scale which is the subject of EFTofLSS [29]. In this section, we will discuss the governing equations for EFTofLSS and the effective stress tensor which comes as result of defining the EFT. We will also discuss significance of this effetcive stress tensor in the context of time non-locality.
@@ 3.1 EFT of LSS
where the subscript "" indicates the long wavelength quantities and is an effective stress tensor generated from the smoothing and contains all the information about short scales. The stress tensor regulates the loops in SPT which access the non-linear scales through integrals in Eq.(2.1) and keeps the theory finite. A generic expansion for the effective stress tensor in terms of the long wavelength dark matter fields is given as,
@@
where are EFT coefficients and is the Kronecker delta function respectively. We can see from Eq.(3.2), that it is local in time as both r.h.s and l.h.s are evaluated at the same time. To write down a local effective theory for the long mode, it is important to note that there should be a separation of length scale. However, for dark matter, even though there is a separation of length scale in the spatial domain but there is no separation of scale in temporal domain. This leads to time non-locality in the effective description of the dark matter after integrating out the short modes. To understand the absence of scale separation in the time domain, we can also look at the linear form of Eq.(2.1). If we linearise the equations given in Eq.(2.1) and substitute the Euler equation in the continuity equation, we get the following linear growth equation for ,
where ′ denotes a derivative w.r.t the conformal time. From Eq.(3.4), we can see that there is no explicit dependence and hence both the long and short modes follow the same dynamics. This is another manifestation of the fact that there is no scale separation in the time domain even though there is a scale separation in the momentum domain through 444 This can also be seen from the discussion above Eq.(2.1), where the dark matter is modelled as a collisionless (pressureless) fluid. Due to the absence of pressure, the speed of sound vanishes in such a fluid. Hence, there does not exist any dispersion relation which relates the wavenumber () with the frequency. However, after integrating out the short modes we get an effective stress tensor on the r.h.s of Eq.(3.4). This introduces an effective pressure and the wavenumber and frequency are related through a dispersion relation.. Therefore, as mentioned earlier, when we smooth out the short modes to write an EFT, the physics is local in space domain but non-locality is introduced in the time domain. This time non-locality is incorporated in the effective stress tensor which can be written as [32, 60],
@@
As we can see from Eq.(3.5), that due to the time integration, Eq.(3.5) is non-local in time as compared to Eq.(3.2) which is local in time. However, we can use Eq.(3.6) iteratively in Eq.(3.5) and Taylor expand it around to get up to any arbitrary order. After expanding about , we will still be left with the integral given in Eq.(3.6). Assuming no vorticity i.e. , we can perform the integral given in Eq.(3.6) and get an expression arranged in powers of [60, 1]. This process removes the time non-locality but introduces space non-locality and the structure of operators appearing in the expansion given in Eq.(3.5) changes [60].
Restricting ourselves upto order in dark matter overdensity, we get [1, 60],
@@ 3.2 Time non-locality in stress tensor ()
We can see from Table. 1, that up to second order the time local and non-local basis are equivalent to each other. However, at third order, the local basis consists of 17 operators while the non-local basis contains 18 independent operators. Hence, we find time non-locality in the effective stress tensor at third order. As a crosscheck of non-local structures, we have computed the independent set of operators in the system as well. We have found that the basis also contains 18 independent operators at third order which is same as the non-local basis. We have checked that both the non-local and bases are related by an invertible map. The list of independent operators is also given in Table. 1.555It’s worth mentioning that despite our focused examination on the non-locality in stress tensor, a similar rationale can be applied to any generic second-rank tensor that the non-locality will shift to third order. Let us note that, at third order, which can be treated as a non-local operator (due to the invertible map), can’t be written in terms of the local operators listed in Table. 1.
@@ 4 Non-locality in biased tracers in real space
Having discussed the local bias expansion in Sec. 2.2, we can discuss time non-locality in biased tracers. In this section, we will layout the physics of biased tracers and discuss the origin of time non-locality among biased tracers. We will mainly be reviewing ideas presented in [54, 60, 1]. We also discuss that writing operators in the system, which is described in Sec. 2.4, makes the task of finding the non-local basis much more efficient.
As mentioned earlier, tracers such as galaxies evolve over the background of dark matter fields. However, it has been shown that the time of virialisation for galaxies or halo is around one-tenth of the time of evolution of dark matter fields i.e. the time taken by a long mode to reach the non-linear scale [54]. This implies that galaxy formation does not occur instantaneously but takes a small but finite time period666If galaxies are formed instantaneously, the background dark matter fields do not change significantly during the formation time. Due to this, the galaxy distribution and hence, the number density will be affected by dark matter fields at the same time. For such a scenario, the bias expansion given in Eq.(2.7) is appropriate as it is local in time in the sense that both r.h.s and l.h.s is evaluated at the same time.. Since, galaxy formation is not instantaneous, the tracers are affected by the dark matter fields not only at the same time but also through the whole past history along the fluid trajectory. Therefore, it has been argued [54], that the actual bias expansion should be non-local in time and the r.h.s of Eq.(2.7) should be generalized to an integral over the fluid trajectory of dark matter as follows,
@@
We can use Eq.(3.6) iteratively to Taylor expand the operator in Eq.(4.1). Restricting ourselves upto order in dark matter overdensity, we get [1, 60],
@@
are the time dependent bias coefficients similar to what is given in the local expansion in Eq.(2.7)777In [1], it has been shown that if the kernels have support over a time scale then the non-local in time bias parameters go like . Therefore, if the bias parameters are order unity, then it implies that the galaxy sample under observation was formed over a timescale of .. From Eq.(4.3), it is clear that are the local in time operators that affect tracer over density at location and time .
@@
Since the virialisation timescale is small as compared to the timescale of evolution of long modes, we expect that to a certain accuracy, the bias expansion can be taken to be local-in-time [54]. However, if we go to sufficiently higher order in overdensity, the local in time and the non-local in time bias expansion must differ in their structure. It has been pointed out [60, 1], that up to fourth order in overdensity, non-local expansion can be written in terms of local expansion i.e the non-local basis and local basis are related by a linear invertible map.
However, at fifth order, the local and the non-local basis are different and hence, there does not exists a linear map between them [1]. It has been shown that at the fifth order, the time-local basis contains 26 independent operators, while the time non-local basis has 29 independent operators. In fact, at this order the total number of non-local operators that appear in Eq.(4.3) are of but the operator space has a lot of degeneracies. After removing degeneracies, the number of independent operators turn out to be only 29 [1]. Out of the 29 non-local operators, 26 of them can be written as linear combinations of the time-local basis and this map is invertible. This implies that the non-local basis contains three extra operators that cannot be written as a linear combinations of operators in the time-local basis. The three non-local terms are pointed out in [1] and are listed below,
@@
In this section, we will show that if we use a different system of writing operators, then the task of finding the complete time non-local basis at any give order becomes swift and easy. In Sec. 2.4, we have described an alternate way of writing operators in the bias expansion [2]. Taking trace over the contraction of ’s and their product, we can construct scalars at any give order. We can now perform the same exercise of writing all possible operators relevant at a given order. Then, we can look for degeneracies among them to get the independent set of operators that will form the basis at that order. We have found that up to fifth order, the basis obtained by -system maps directly to the non-local basis through an invertible linear map. At fifth order, we found that there are a total of 52 scalar operators out of which only 29 are independent. The list of 29 independent operators at order five is given as follows,
where we have arranged the set of operators by the order at which they start and we are denoting it by index m. Also we are using the notation . As mentioned earlier, the basis that we get from the system given in Eq.(4.6) and from the non-local expansion which is given in Eq.(4.3) are related by an invertible map up to fifth order. We have shown the explicit mapping in the Mathematica file attached with the manuscript. Here we will just write the three "non-local" operators as pointed in [1] as a linear combination of operators,
As mentioned earlier (Eq.(4.5)), the specific non-local operators cannot be written in terms of the time-local basis but we have shown in Eq.(4.7), they can indeed be written as a linear combination of operators given in the alternative basis, Eq.(4.6). We would like to point out that the system is more efficient in getting the non-local basis as we just get 52 number of scalar operators which is much less in contrast to around 130 scalar operators we get from the usual Taylor expansion in Eq.(4.3). As a result, at fifth order, it has less number of degenerate operators888Only 23 degenerate operators are present at fifth order in the system. as compared to the non-local expansion which contains degenerate operators and hence a lot of redundancies. This reduces the task of finding degeneracy equations which is one of the most challenging aspect of this calculation. Note that Eq.(4.6) does not contain operators involving velocity gradient. We have checked that the operators constructed from velocity gradient and it’s convective derivatives are degenerate with the basis given in Eq.(4.6) up to fifth order. This was pointed out in [2].
@@ 5.1 Non-locality shifts to order
As mentioned earlier, selection operators are needed to take into the line of sight dependent effects that can affect the detection of tracer overdensity. In this section, we discuss the set of independent selection operators appearing at third order. Note that the numbers we quote for independent operators is specific to selection operators only. The full bias expansion, however, will consist of both the as well as the selection operators. For the case of Lyman- forest, all such operators are generated by loops through the exponential map even if they are not present at tree level [57]. So we need to include all possible operators with the structure given in Eq.(2.15). Constructing all possible contractions we get the set of local operators. After removing the degenerate operators we find that at third order in tracer overdensity, there are 15 independent local selection operators999A subset of the operators shown in Eq.(5.1) are generated when writing the bias expansion in redshift space. However, we have checked that up to third order, the local and the non-local basis are equivalent.. The basis of selection operators at third order is given below,
@@
We can also write selection operators in the system used in Sec.2.4. For that we need to make two indexed objects from defined in Eq.(2.17) and take projection with the unit vectors along the line of sight such as . At third order, we find only 16 independent operators. The list of independent selection operators up to third order in dark matter over density is as follows,
where we have arranged the operators by the index “” that denotes the order in dark matter overdensity at which a particular operator starts. As was the case for operators, we find that here also there exists an invertible map between the non-local basis given in Eq.(5.2) and basis as given in Eq.(5.3). We have shown the explicit mapping in the Mathematica file attached with the manuscript. We can see from Eq.(5.3), that it does not contain operators involving velocity gradient. This is due to the fact that the operators constructed from velocity gradient () and its convective derivatives are degenerate with the basis given in Eq.(5.3). This was pointed out in [2, 61] and we have explicitly verified this.
The relation between non-local expansion and the basis holds true at fourth order as well. We will explore this in a bit more detail ahead. As for the case of third order, we can write the time-local selection operators up to fourth order by constructing scalars. We construct 67 such operators out of which only 41 are independent and form the basis at fourth order. The independent operators are listed below,
where we have followed the notation given in Eq.(2.9) and Eq.(2.15) for writing operators in a compact form. To get the time non-local basis we need to Taylor expand the operators in Eq.(5.1) along the fluid trajectory as we did for the case of operators given in Eq.(4.3). At fourth order, we have found 133 such non-local operators, out of which 87 are degenerate and therefore only 46 operators are independent which we list out here,
@@
So Eq.(5.6) together gives the total number of selection operators at fourth order. From Eq.(5.6) we can see that there are 30 new operators that appear at fourth order coming from sector, which makes the total number of selection operators at order four to be 46.
The operators give in Eq.(5.6) are related to the time non-local basis with an invertible map. We have explicitly shown the mapping in the Mathematica file attached with the manuscript.101010Note that Eq.(5.6) does not contain operators with velocity gradient () and it’s convective derivatives. This is because at fourth order also, all such operators are degenerate with the operators given in Eq.(5.6). At fourth order, we can write 53 operators involving velocity gradient and their convective derivatives. We have explicitly checked that each of them is degenerate with the operators given in Eq.(5.6).

### v2

@@ 3 Non-locality in dark matter
SPT has been sucessful in describing LSS at very large scales but as we approach shorter distances i.e. higher wavenumber, linear SPT starts to diverge from the results of fully nonlinear -body simulations. Even for , i.e. quasilinear scales where , higher loop calculations of the SPT power spectrum do not converge to the correct result [Carrasco:2013mua]. To cure this, the correct treatment is to smooth out the equations of motion over a length scale above the non-linear scale. The effect of this smoothing out procedure is to include “effective" contributions from non-linear scale which is the subject of EFTofLSS [Carrasco:2012cv]. In this section, we will discuss the governing equations for EFTofLSS and the effective stress tensor which comes as result of defining the EFT. We will also discuss significance of this effetcive stress tensor in the context of time non-locality.
@@ 3.1 EFT of LSS
where the subscript "" indicates the long wavelength quantities and is an effective stress tensor generated from the smoothing and contains all the information about short scales. The stress tensor regulates the loops in SPT which access the non-linear scales through integrals in Eq.(2.1) and keeps the theory finite. One can write the effective stress tensor in terms of the long wavelength dark matter fields as follows,
@@
where are EFT coefficients and is the Kronecker delta function respectively. We can see from Eq.(3.2), that it is local in time as both r.h.s and l.h.s are evaluated at the same time. Note that Eq.(3.2) is not the most general expression for the effective stress tensor. The stress tensor is actually non-local in time for reasons we describe below. Therefore, the most general expression for the effective stress tensor is also non-local in time and is given later in Eq.(3.5). To write down a local effective theory for the long mode, it is important to note that there should be a separation of length scale. However, for dark matter, even though there is a separation of length scale in the spatial domain but there is no separation of scale in temporal domain. This leads to time non-locality in the effective description of the dark matter after integrating out the short modes. To understand the absence of scale separation in the time domain, we can also look at the linear form of Eq.(2.1). If we linearise the equations given in Eq.(2.1) and substitute the Euler equation in the continuity equation, we get the following linear growth equation for ,
where ′ denotes a derivative w.r.t the conformal time. From Eq.(3.4), we can see that there is no explicit dependence and hence both the long and short modes follow the same dynamics. This is another manifestation of the fact that there is no scale separation in the time domain even though there is a scale separation in the momentum domain through 444This can also be seen from the discussion above Eq.(2.1), where the dark matter is modelled as a collisionless (pressureless) fluid. Due to the absence of pressure, the speed of sound vanishes in such a fluid. Hence, there does not exist any dispersion relation which relates the wavenumber () with the frequency. However, after integrating out the short modes we get an effective stress tensor on the r.h.s of Eq.(3.4). This introduces an effective pressure and the wavenumber and frequency are related through a dispersion relation.. Therefore, as mentioned earlier, when we smooth out the short modes to write an EFT, the physics is local in space domain but non-locality is introduced in the time domain. The time non-locality in the context of EFTofLSS for dark matter was first pointed in [Carrasco:2013mua]. The time non-locality is incorporated in the effective stress tensor which can be written as [Carrasco:2013mua, DAmico:2022ukl],
@@
As we can see from Eq.(3.5), that due to the time integration, Eq.(3.5) is non-local in time as compared to Eq.(3.2) which is local in time. However, we can use Eq.(3.6) iteratively in Eq.(3.5) and Taylor expand it around to get up to any arbitrary order. After expanding about , we will still be left with the integral given in Eq.(3.6). Assuming no vorticity i.e. , we can perform the integral given in Eq.(3.6) and get an expression arranged in powers of [DAmico:2022ukl, Donath:2023sav]. This process removes the time non-locality but introduces space non-locality and the structure of operators appearing in the expansion given in Eq.(3.5) changes [DAmico:2022ukl].
Restricting ourselves upto order in dark matter overdensity, we get [Donath:2023sav, DAmico:2022ukl],
@@
Note that Taylor expanding Eq.(3.5) using Eq.(3.6) introduces the convective derivative structure similar to what is given in Eq.(2.17). Therefore, one expects that the basis obtained in this manner is generically equivalent to the basis discussed in Sec. 2.4, which is constructed by acting convective derivatives on local operators. We will see later that indeed both systems produce equivalent bases at each order.
@@ 3.2 Time non-locality in stress tensor ()
We can see from Table. 1, that up to second order the time local and non-local basis are equivalent to each other. However, at third order, the local basis consists of 16 operators while the non-local basis contains 17 independent operators. Hence, we find time non-locality in the effective stress tensor at third order. As a crosscheck of non-local structures, we have computed the independent set of operators in the system as well. Note that the general structure of effective stress tensor has been obtained up to third order in terms of basis elsewhere [Vlah:2019byq]555We thank the referee for pointing this out to us.. We have found that the basis also contains 17 independent operators at third order which is same as the non-local basis. Our result for basis agrees with that obtained in [Vlah:2019byq]666The list of operators given in [Vlah:2019byq] contains 11 tensor structures. However, it can be shown that one of the operators can be written as a linear combination of other tensor structures and times trace operators. This makes the list of independent tensor structures at third order to be 10 which is given in Table.1.. We have checked that both the non-local and bases are related by an invertible map. The list of independent operators is also given in Table. 1.777It’s worth mentioning that despite our focused examination on the non-locality in stress tensor, a similar rationale can be applied to any generic second-rank tensor that the non-locality will shift to third order. Let us note that, at third order, which can be treated as a non-local operator (due to the invertible map), can’t be written in terms of the local operators listed in Table. 1888We note that our list of operators included in the trace part of contains whereas the corresponding list in [Vlah:2019byq] features . We would like to emphasise that both are equivalent choices at third order..
@@ 4 Non-locality in biased tracers in real space
Having discussed the local bias expansion in Sec. 2.2, we can discuss time non-locality in biased tracers. In this section, we will layout the physics of biased tracers and discuss the origin of time non-locality among biased tracers. We will mainly be reviewing ideas presented in [Senatore:2014eva, DAmico:2022ukl, Donath:2023sav]. We also discuss that writing operators in the system, which is described in Sec. 2.4, makes the task of finding the non-local basis much more efficient.
As mentioned earlier, tracers such as galaxies evolve over the background of dark matter fields. However, it has been shown that the time of virialisation for galaxies or halos is around one-tenth of the time of evolution of dark matter fields i.e. the time taken by a long mode to reach the non-linear scale [Senatore:2014eva]. This implies that galaxy formation does not occur instantaneously but takes a small but finite time period999If galaxies are formed instantaneously, the background dark matter fields do not change significantly during the formation time. Due to this, the galaxy distribution and hence, the number density will be affected by dark matter fields at the same time. For such a scenario, the bias expansion given in Eq.(2.6) is appropriate as it is local in time in the sense that both r.h.s and l.h.s is evaluated at the same time.. Since, galaxy formation is not instantaneous, the tracers are affected by the dark matter fields not only at the same time but also through the whole past history along the fluid trajectory. Therefore, it has been argued [Senatore:2014eva], that the actual bias expansion should be non-local in time and the r.h.s of Eq.(2.6) should be generalized to an integral over the fluid trajectory of dark matter as follows,
@@
Following the same procedure as we did in Sec.3.1, we can use Eq.(3.6) iteratively to Taylor expand the operator in Eq.(4.1). Restricting ourselves upto order in dark matter overdensity, we get [Donath:2023sav, DAmico:2022ukl],
@@
are the time dependent bias coefficients similar to what is given in the local expansion in Eq.(2.6)101010In [Donath:2023sav], it has been shown that if the kernels have support over a time scale then the non-local in time bias parameters go like . Therefore, if the bias parameters are order unity, then it implies that the galaxy sample under observation was formed over a timescale of .. From Eq.(4.3), it is clear that are the local in time operators that affect tracer over density at location and time .
@@
Since the virialisation timescale is small as compared to the timescale of evolution of long modes, we expect that to a certain accuracy, the bias expansion can be taken to be local-in-time [Senatore:2014eva]. However, if we go to sufficiently higher order in overdensity, the local in time and the non-local in time bias expansion must differ in their structure. It has been pointed out [DAmico:2022ukl, Donath:2023sav], that up to fourth order in overdensity, non-local expansion can be written in terms of local expansion i.e the non-local basis and local basis are related by a linear invertible map.
However, at fifth order, the local and the non-local basis are different and hence, there does not exists a linear map between them [Donath:2023sav]. It has been shown that at the fifth order, the time-local basis contains 26 independent operators, while the time non-local basis has 29 independent operators. In fact, at this order the total number of non-local operators that appear in Eq.(4.3) are of but the operator space has a lot of degeneracies. After removing degeneracies, the number of independent operators turn out to be only 29 [Donath:2023sav]. Out of the 29 non-local operators, 26 of them can be written as linear combinations of the time-local basis and this map is invertible. This implies that the non-local basis contains three extra operators that cannot be written as a linear combinations of operators in the time-local basis. The three non-local terms are pointed out in [Donath:2023sav] and are listed below,
@@
In this section, we will show that if we use a different system of writing operators, then the task of finding the complete time non-local basis at any give order becomes swift and easy. In Sec. 2.4, we have described an alternate way of writing operators in the bias expansion [Mirbabayi:2014zca]. Taking trace over the contraction of ’s and their product, we can construct scalars at any give order. We can now perform the same exercise of writing all possible operators relevant at a given order. Then, we can look for degeneracies among them to get the independent set of operators that will form the basis at that order. We have found that up to fifth order, the basis obtained by -system maps directly to the non-local basis through an invertible linear map. At fifth order, we found that there are a total of 52 scalar operators out of which only 29 are independent. The list of 29 independent operators at order five is given as follows,
where we have arranged the set of operators by the order at which they start and we are denoting it by index m. Also we are using the notation . As mentioned earlier, the basis that we get from the system given in Eq.(4.6) and from the non-local expansion which is given in Eq.(4.3) are related by an invertible map up to fifth order. We have found the explicit mapping using Mathematica111111The Mathematica files are attached with the preprint version at arxiv: 2406.17025.. Here we will just write the three "non-local" operators as pointed in [Donath:2023sav] as a linear combination of operators,
As mentioned earlier (Eq.(4.5)), the specific non-local operators cannot be written in terms of the time-local basis but we have shown in Eq.(4.7), they can indeed be written as a linear combination of operators given in the alternative basis, Eq.(4.6). We would like to point out that the system is more efficient in getting the non-local basis as we just get 52 number of scalar operators which is much less in contrast to around 130 scalar operators we get from the usual Taylor expansion in Eq.(4.3). As a result, at fifth order, it has less number of degenerate operators121212Only 23 degenerate operators are present at fifth order in the system. as compared to the non-local expansion which contains degenerate operators and hence a lot of redundancies. This reduces the task of finding degeneracy equations which is one of the most challenging aspect of this calculation. Note that Eq.(4.6) does not contain operators involving velocity gradient. We have checked that the operators constructed from velocity gradient and it’s convective derivatives are degenerate with the basis given in Eq.(4.6) up to fifth order. This was pointed out in [Mirbabayi:2014zca].
@@ 5.1 Non-locality shifts to order
As mentioned earlier, selection operators are needed to take into the line of sight dependent effects that can affect the detection of tracer overdensity. In this section, we discuss the set of independent selection operators appearing at third order. Note that the numbers we quote for independent operators is specific to selection operators only. The full bias expansion, however, will consist of both the as well as the selection operators. For the case of Lyman- forest, all such operators are generated by loops through the exponential map even if they are not present at tree level [Ivanov:2023yla]. So we need to include all possible operators with the structure given in Eq.(2.14). Constructing all possible contractions we get the set of local operators. After removing the degenerate operators we find that at third order in tracer overdensity, there are 15 independent local selection operators131313A subset of the operators shown in Eq.(5.1) are generated when writing the bias expansion in redshift space. However, we have checked that up to third order, the local and the non-local basis in redshift space are equivalent, which has also been noted elsewhere [DAmico:2022ukl].. The basis of selection operators at third order is given below,
@@
We can also write selection operators in the system used in Sec.2.4. For that one needs to make two indexed objects from defined in Eq.(2.16) and take projection with the unit vectors along the line of sight such as . At the third order, one finds only 16 independent operators. The list of independent selection operators up to third order in dark matter overdensity is as follows [Desjacques:2018pfv],
where we have arranged the operators by the index “” that denotes the order in dark matter overdensity at which a particular operator starts. As was the case for operators, we find that here also there exists an invertible map between the non-local basis given in Eq.(5.2) and basis as given in Eq.(5.3). We have found the invertible mapping explicitly using Mathematica. We can see from Eq.(5.3), that it does not contain operators involving velocity gradient. This is due to the fact that the operators constructed from velocity gradient () and its convective derivatives are degenerate with the basis given in Eq.(5.3). This was pointed out in [Mirbabayi:2014zca, Zaldarriaga:2015jrj] and we have explicitly verified this.
The relation between non-local expansion and the basis holds true at fourth order as well. We will explore this in a bit more detail ahead. As for the case of third order, we can write the time-local selection operators up to fourth order by constructing scalars. We construct 67 such operators out of which only 40 are independent and form the basis at fourth order. The independent operators are listed below,
where we have followed the notation given in Eq.(2.8) and Eq.(2.14) for writing operators in a compact form. To get the time non-local basis we need to Taylor expand the operators in Eq.(5.1) along the fluid trajectory as we did for the case of operators given in Eq.(4.3). At fourth order, we have found 133 such non-local operators, out of which 88 are degenerate and therefore only 45 operators are independent which we list out here,
@@
So Eq.(5.6) together gives the total number of selection operators at fourth order. From Eq.(5.6) we can see that there are 29 new operators that appear at fourth order coming from sector, which makes the total number of selection operators at order four to be 45.
The operators given in Eq.(5.6) are related to the time non-local basis with an invertible map. We have found the explicit mapping using Mathematica.141414Note that Eq.(5.6) does not contain operators with velocity gradient () and it’s convective derivatives. This is because at fourth order also, all such operators are degenerate with the operators given in Eq.(5.6). At fourth order, we can write 53 operators involving velocity gradient and their convective derivatives. We have explicitly checked that each of them is degenerate with the operators given in Eq.(5.6).